# Deep Generative Methods and Tire Architecture Design

### v1

@@ Abstract
As deep generative models proliferate across the AI landscape, industrial practitioners still face critical yet unanswered questions about which deep generative models best suit complex manufacturing design tasks. This work addresses this question through a complete study of five representative models (Variational Autoencoder, Generative Adversarial Network, multimodal Variational Autoencoder, Denoising Diffusion Probabilistic Model, and Multinomial Diffusion Model) on industrial tire architecture generation. Our evaluation spans three key industrial scenarios: (i) unconditional generation of complete multi-component designs, (ii) component-conditioned generation (reconstructing architectures from partial observations), and (iii) dimension-constrained generation (creating designs that satisfy specific dimensional requirements). To enable discrete diffusion models to handle conditional scenarios, we introduce categorical inpainting, a mask-aware reverse diffusion process that preserves known labels without requiring additional training. Our evaluation employs geometry-aware metrics specifically calibrated for industrial requirements, quantifying spatial coherence, component interaction, structural connectivity, and perceptual fidelity. Our findings reveal that diffusion models achieve the strongest overall performance; a masking-trained VAE nonetheless outperforms the multimodal variant MMVAE+ on nearly all component-conditioned metrics, and within the diffusion family MDM leads in-distribution whereas DDPM generalises better to out-of-distribution dimensional constraints.
@@ II-E Benchmarking Generative Models for Industrial Applications
To the best of our knowledge, systematic benchmarks that compare multiple DGMs under identical conditions are still rare in the industrial‐design literature, yet a few recent works offer useful precedents. In materials science, [28] evaluated six generative architectures—including VAEs, Adversarial Autoencoders, several GAN variants, and an RNN baseline—on a common polymer dataset, reporting validity, novelty, and property‐matching scores. Mechanical‐metamaterial research has likewise seen a four-way comparison (VAE, GAN, WGAN, DDPM) on kirigami cut-pattern generation, which highlighted the difficulty of satisfying hard geometric constraints with Euclidean latent spaces [29]. For large disordered materials, the Dismai-Bench framework contrasted graph-based and coordinate-based diffusion models (plus point-cloud GANs), demonstrating the importance of symmetry-aware data representations [30]. Outside design, vibration-based fault diagnosis in rotating machinery has been benchmarked with convolutional and variational autoencoders, revealing trade-offs between reconstruction fidelity and anomaly sensitivity [31]. Collectively, these works establish that no single family of DGMs dominates across domains and metrics.
@@ III-A Notation
  • Categorical representation (MDM): for each pixel $i\in\{1,\dots,H\!\times\!W\}$ we define $\mathbf{c}_{i}\in\{0,1\}^{K}$ as its one-hot label ($K=M{+}1$, i.e. $M$ components plus background). For simplicity we omit the pixel index.
@@
  • $\mathbf{d}=(w,h)$ represents the dimensional parameters of a tire architecture, where $w$ is the width and $h$ is the height. For dimension-constrained generation, $\mathbf{d}_{\text{target}}=(w_{\text{target}},h_{\text{target}})$ denotes the target dimensions, and $\mathbf{d}_{\text{actual}}=(w_{\text{actual}},h_{\text{actual}})$ represents the actual dimensions of a generated sample.
@@ III-B Dataset and Model-Specific Preprocessing
[…]
@@ III-C5 Multinomial Diffusion Model (MDM)
where $K=M{+}1$ and $\mathbf{u}$ is the uniform probability vector.
@@
Categorical inpainting.
Analogous to the Gaussian trick of [32], we multiply every reverse step by a binary mask that freezes the observed labels. Let $\mathbf{c}_{0}\!\in\!\{0,1\}^{K\times H\times W}$ and $\mathcal{O}$ be the set of observed pixels. For each pixel we set $m_{k}=\mathbf{1}_{\{c_{0}^{i}\}}(k)$ if $i\!\in\!\mathcal{O}$ and $m_{k}\!=\!1$ otherwise.
@@
Let $\mathbf{r}=(r_{0},\ldots,r_{K})$ be any non-negative weight vector with $\sum_{k}r_{k}\theta_{k}>0$.
@@
$\tilde{p}(\mathbf{c}_{t-1}|\mathbf{c}_{t})=\frac{\mathbf{r(\mathbf{c}_{t-1})}\cdot p(\mathbf{c}_{t-1}|\mathbf{c}_{t})}{\sum_{j=0}^{K}r_{j}\cdot\theta_{j}}$ (9)

### v2

@@ Abstract
As deep generative models proliferate across the AI landscape, industrial practitioners still face critical yet unanswered questions about which deep generative models best suit complex manufacturing design tasks. This work addresses this question through a complete study of five representative models (Variational Autoencoder, Generative Adversarial Network, multimodal Variational Autoencoder, Denoising Diffusion Probabilistic Model, and Multinomial Diffusion Model) on industrial tire architecture generation. Our evaluation spans three key industrial scenarios: (i) unconditional generation of complete multi-component designs, (ii) component-conditioned generation (reconstructing architectures from partial observations), and (iii) dimension-constrained generation (creating designs that satisfy specific dimensional requirements). To enable discrete diffusion models to handle conditional scenarios, we introduce categorical inpainting, a mask-aware reverse diffusion process that preserves known labels without requiring additional training. Our evaluation employs geometry-aware metrics specifically calibrated for industrial requirements, quantifying spatial coherence, component interaction, structural connectivity, and perceptual fidelity. Our findings reveal that diffusion models achieve the strongest overall performance; a masking-trained VAE nonetheless outperforms the multimodal variant MMVAE+ on nearly all component-conditioned metrics, and within the diffusion family MDM leads in-distribution whereas DDPM generalises better to out-of-distribution dimensional constraints.  [^1] An anonymized implementation is available at https://anonymous.4open.science/r/DGMTA-5122/. A fully documented, open‑source repository will be released upon acceptance.
@@ II-E Benchmarking Generative Models for Industrial Applications
To the best of our knowledge, systematic benchmarks that compare multiple DGMs under identical conditions are still rare in the industrial‐design literature, yet a few recent works offer useful precedents. In materials science, [28] evaluated six generative architectures (VAEs, Adversarial Autoencoders, several GAN variants, and an RNN baseline), on a common polymer dataset, reporting validity, novelty, and property‐matching scores. Mechanical‐metamaterial research has likewise seen a four-way comparison (VAE, GAN, WGAN, DDPM) on kirigami cut-pattern generation, which highlighted the difficulty of satisfying hard geometric constraints with Euclidean latent spaces [29]. For large disordered materials, the Dismai-Bench framework contrasted graph-based and coordinate-based diffusion models (plus point-cloud GANs), demonstrating the importance of symmetry-aware data representations [30]. Outside design, vibration-based fault diagnosis in rotating machinery has been benchmarked with convolutional and variational autoencoders, revealing trade-offs between reconstruction fidelity and anomaly sensitivity [31]. Collectively, these works establish that no single family of DGMs dominates across domains and metrics.
@@ III-A Notation
  • Categorical representation (MDM): for each pixel $i\in\{1,\dots,H\!\times\!W\}$ we define $\mathbf{c}_{i}\in\{0,1\}^{K}$ as its one-hot label (K=M+1K=M{+}1, i.e. $M$ components plus background). For simplicity we write $\mathbf{c}$ instead of $\mathbf{c}_{i}$ whenever the index is unambiguous.
@@
  • $\mathbf{d}=(w,h)$ represents the dimensional parameters of a tire architecture, where $w$ is the width and $h$ is the height.
@@ III-B Dataset and Model-Specific Preprocessing
Our proprietary dataset consists of 1,130 tire architectures, each composed of six distinct structural components: $\{\mathit{bead\ filler},\mathit{cushion},\mathit{carcass},\mathit{sidewall},\mathit{tread},\mathit{inner\ liner}\}$ (Figure 1). More industrial details about these components and their functional roles can be found in the appendix. Each component is represented as a grayscale image of resolution $64\times 64$. We partition the dataset into 80% for training and 20% for testing.
@@ III-C5 Multinomial Diffusion Model (MDM)
where K=M+1K=M{+}1 and $\mathbf{u}$ is the uniform probability vector.
@@
Categorical inpainting
Analogous to the Gaussian “delta–mask” trick of [32], we multiply every reverse step by a binary mask that freezes the observed labels. Let $\mathbf{c}_{0}\!\in\!\{0,1\}^{K\times H\times W}$ and $\mathcal{O}$ be the set of observed pixels. For each pixel we set the mask $r_{k}=(\mathbf{c}_{0}^{i})_{k}$, $k\in\{1,\dots,K\}$ if $i\in\mathcal{O}$ and $r_{k}=1$ otherwise.
@@
Let $\mathbf{r}=(r_{1},\ldots,r_{K})$ be any non-negative weight vector with $\sum_{k}r_{k}\theta_{k}>0$.
@@
$\tilde{p}(\mathbf{c}_{t-1}|\mathbf{c}_{t})=\frac{\mathbf{r(\mathbf{c}_{t-1})}\cdot p(\mathbf{c}_{t-1}|\mathbf{c}_{t})}{\sum_{j=1}^{K}r_{j}\cdot\theta_{j}}$ (9)
@@
Where $\mathbf{r}(\mathbf{c}_{t-1})$ denotes the scalar product $\langle\mathbf{r},\mathbf{c}_{t-1}\rangle$.